OpenStudy (anonymous):

What is the remainder of (2x3 − 18x + 5) ÷ (x − 3)? -5 5 2 3

OpenStudy (anonymous):

@Peaches15 can you help?

OpenStudy (anonymous):

4

Nnesha (nnesha):

x-5=0 solve for x and the substitute x in this 2x^3 − 18x + 5 equation for its value

Nnesha (nnesha):

or you can apply synthetic division

OpenStudy (sohailiftikhar):

it should be -5

OpenStudy (anonymous):

-5 is wrong i got it wrong the last time i tried it

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha x-5=0 solve for x and the substitute x in this 2x^3 − 18x + 5 equation for its value \(\color{blue}{\text{End of Quote}}\) correction: x-3=0

OpenStudy (anonymous):

would it be 3?

OpenStudy (sohailiftikhar):

ok clarify your equation 2x3 or 2x^3

OpenStudy (sohailiftikhar):

solve for this x-3=0 then x=3 put value of x in equation you will get answer

OpenStudy (anonymous):

sorry its 2x^3

Nnesha (nnesha):

yep so now substitute al x for 3\[\huge\rm 2(3)^3-18(3)+5= ??\]

OpenStudy (sohailiftikhar):

it's 5

OpenStudy (anonymous):

so 2-18+5=-11

OpenStudy (sohailiftikhar):

lol nah put value of x

OpenStudy (sohailiftikhar):

you will get 54-54+5=5

Nnesha (nnesha):

no substitute x for 3 just replace x

OpenStudy (anonymous):

oh okay i think i get it now

Nnesha (nnesha):

x for 3\[\huge\rm 2(\color{reD}{3})^3-18(\color{reD}{3})+5= ??\]

OpenStudy (sohailiftikhar):

when you are given like that divided by x-5 or x-3 etc then just put it equal to zero and get value of x put this value in equation you will get your answer

OpenStudy (sohailiftikhar):

good job Nnesha

Nnesha (nnesha):

or if you want to apply synthetic division then you need to write highest degree to smallest  0 for 0x^2 it's not in the equation but we still have to write highest degree to smallest x^3,x^2 and then x